# Connectedness of fibers beyond semitoric systems I: the non-degenerate case

### v3

@@ 7. Proof of the main result
In this section, we prove of the main results, Theorems 1.6 and 1.10. To begin, we explain how certain questions about the topology of the integrable systems we consider can be reduced to understanding Morse theory for oriented surfaces.
@@
If $\Phi^{-1}(\beta_{0})/T$ is simply connected for some $\beta_{0}\in\Phi(M)$, then M//T{M/\!/T} is simply connected by [KT1, Lemma 5.7 and Corollary 9.7]. Therefore, the fibers $f^{-1}(\beta,c)$ are connected for all $c\in{\mathbb{R}}$ exactly if M//T{M/\!/T} has no tall singular points with a hyperbolic block and connected $T$-stabilizer by Lemma 7.3.
@@
Let $N$ be a connected manifold and let $h\colon N\to{\mathbb{R}}$ be a proper function such that $h(N)\subseteq[0,\infty)$. Assume that the restriction of $h$ to $h^{-1}(0,\infty)$ is a Morse-Bott function with no points of index $1$, and that 0 is an isolated critical value. Then $h^{-1}(0)$ is connected.
@@
Consider points $x$ and $y$ in $h^{-1}(0)$. Since $N$ is path-connected, there is a path $\gamma$ from $x$ to $y$. Hence, $x$ and $y$ are in the same component of $h^{-1}([0,b))$ for some $b\in{\mathbb{R}}$. If $C\subseteq h^{-1}((0,\infty))$ is a critical submanifold, then since the index of $C$ is not $1$, the restriction $H^{0}(h^{-1}([0,f(C)+\epsilon)))\to H^{0}(h^{-1}([0,f(C)-\epsilon)))$ is a surjection for sufficiently small $\epsilon>0$. In particular, if $x$ and $y$ are in the same component of $h^{-1}([0,h(C)+\epsilon))$, then they are also in the same component of $h^{-1}([0,h(C)-\epsilon))$. Moreover, since $h$ is proper and 0 is an isolated critical value, there are a finite number of such critical submanifolds in $h^{-1}([0,b))$. Hence, by induction, $x$ and $y$ are in the same component of $h^{-1}(0).$
@@
Fix $c\in{\mathbb{R}}$ and define $h\colon N\to{\mathbb{R}}$ by $h=(\overline{g}-c)^{2}$. On the set $\overline{g}^{-1}(c,\infty)$ the function $h$ is Morse-Bott; the critical components of $h$ and $\overline{g}$ coincide and have the same index. On the set $\overline{g}^{-1}(-\infty,c)$ the function $h$ is Morse-Bott; the critical components of $h$ and $\overline{g}$ coincide but the index and coindex of each critical manifold are exchanged. Finally, since $h$ is a proper Morse-Bott function there exists a neighborhood of $c$ that contains no other critical values, so 0 is an isolated critical value. Therefore, $h^{-1}(0)=\overline{g}^{-1}(c)$ is connected by Lemma 7.5. ∎

### v4

@@ 7. Proof of the main result
In this section, we prove of the main results, Theorems 1.6 and 1.11. To begin, we explain how certain questions about the topology of the integrable systems we consider can be reduced to understanding Morse theory for oriented surfaces.
@@
If $\Phi^{-1}(\beta_{0})/T$ is simply connected for some $\beta_{0}\in\Phi(M)$, then $\Phi^{-1}(\beta^{\prime})/T$ is simply connected for all $\beta^{\prime}\in\Phi(M)$ by [KT1, Lemma 5.7 and Corollary 9.7]. In particular, M//T{M/\!/T} is simply connected. Therefore, the fibers $f^{-1}(\beta,c)$ are connected for all $c\in{\mathbb{R}}$ exactly if M//T{M/\!/T} has no tall singular points with a hyperbolic block and connected $T$-stabilizer by Lemma 7.3.
@@
The following result partially generalizes Theorem 1.6.
Proposition 7.5.
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g)\right)$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space with connected fibers. Assume that $f$ is proper, that each tall singular point in $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$ is non-degenerate, and that no such point has both a hyperbolic block and connected $T$-stabilizer. Then the fiber $f^{-1}(\beta,c)$ is connected for each $(\beta,c)\in{\mathfrak{t}}^{*}\times{\mathbb{R}}$.
Proof.
Fix $\beta\in{\mathfrak{t}}^{*}$ such that $\Phi^{-1}(\beta)/T$ contains more than one point; we may assume that $\beta=0$. By assumption, the reduced space M//T{M/\!/T} is connected and $\overline{g}\colon{M/\!/T}\to{\mathbb{R}}$ is proper. Hence, every point in M//T{M/\!/T} is tall, and so by Proposition 7.1, M//T{M/\!/T} is a smooth, connected oriented surface so that $\overline{g}\colon{M/\!/T}\to{\mathbb{R}}$ is a Morse function; moreover, $\overline{g}$ has no critical points of index $1$. Therefore, by Proposition 5.6 the fiber $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$. Finally, since $T$ is connected, this implies that the fiber $f^{-1}(0,c)$ is connected for all $c\in{\mathbb{R}}$. ∎
@@
Let $N$ be a connected manifold and let $h\colon N\to{\mathbb{R}}$ be a proper function such that $h(N)\subseteq[0,\infty)$. Assume that the restriction of $h$ to $h^{-1}(0,\infty)$ is a Morse-Bott function with no points of index $1$. Then $h^{-1}(0)$ is connected.
@@
Consider points $x$ and $y$ in $h^{-1}(0)$. Since $N$ is path-connected, there is a path $\gamma$ from $x$ to $y$. Hence, $x$ and $y$ are in the same component of $h^{-1}([0,b))$ for some $b\in{\mathbb{R}}$. If $C\subseteq h^{-1}((0,\infty))$ is a critical submanifold, then since the index of $C$ is not $1$, the restriction $H^{0}(h^{-1}([0,f(C)+\epsilon)))\to H^{0}(h^{-1}([0,f(C)-\epsilon)))$ is a surjection for sufficiently small $\epsilon>0$. In particular, if $x$ and $y$ are in the same component of $h^{-1}([0,h(C)+\epsilon))$, then they are also in the same component of $h^{-1}([0,h(C)-\epsilon))$. Moreover, since $h$ is proper and its restriction to $h^{-1}(0,\infty)$ is Morse-Bott, the set of critical values in $(\delta,b)$ is finite for all regular $\delta\in(0,b)$. By induction on these critical values, $x$ and $y$ lie in the same connected component $X_{\delta}$ of $h^{-1}([0,\delta])$ for all $\delta\in(0,b)$. Since $N$ is Hausdorff, and $X_{\delta}\subseteq X_{\delta^{\prime}}$ are compact and connected for all $\delta^{\prime}<\delta$, the intersection $\cap_{\delta}X_{\delta}$ is connected. ∎
@@
Let $N$ be a connected manifold and let $\overline{g}\colon N\to{\mathbb{R}}$ be a proper Morse-Bott function. Fix $c\in{\mathbb{R}}$ and consider the proper function $h\colon N\to{\mathbb{R}}$ given by $h=(\overline{g}-c)^{2}$. On the set $\overline{g}^{-1}(c,\infty)$ the function $h$ is Morse-Bott; the critical components of $h$ and $\overline{g}$ coincide and have the same index. On the set $\overline{g}^{-1}(-\infty,c)$ the function $h$ is Morse-Bott; the critical components of $h$ and $\overline{g}$ coincide but the index and coindex of each critical manifold are exchanged. Therefore, if no critical point of $\overline{g}$ has index or coindex one, then $\overline{g}^{-1}(c)=h^{-1}(0)$ is connected by Lemma 7.6. ∎